OpenStudy (13ziar):

In solving equations how do you know weather to add or subtract. eg 6x+6y=24 -5x+6y=21

OpenStudy (anonymous):

If the number you're trying to neutralize is positive, you subtract and if negative you add.

OpenStudy (anonymous):

inverse of each operation...

OpenStudy (13ziar):

I've been taught that both equations add so you do the inverse so for that equation you would subtract the top from the bottom

OpenStudy (anonymous):

are solving solving both equation at the same time? or individually?

OpenStudy (13ziar):

at the same time

OpenStudy (anonymous):

then use elimination method

OpenStudy (13ziar):

yes thats what im doing then substituting it

OpenStudy (13ziar):

what about this equation: 6x+7y=-16 2x-4y=5

OpenStudy (anonymous):

ok the idea is to cancel one variable, and solve the other variable

OpenStudy (anonymous):

ok what variable u want to cancel first..

OpenStudy (anonymous):

choose wisely

OpenStudy (anonymous):

6x+7y=-16 2x-4y=5 if you multiply the second equation by -3, ull get -6x+12y=15

OpenStudy (13ziar):

OMG i was typing and wrote and it just started to reload the page

OpenStudy (13ziar):

ok so ... basically i know how to solve them but ....

OpenStudy (13ziar):

.... i want to know that if you have this equation: 6x+7y=-16 2x-4y=5 then to eliminate/solve how do you know if you add or subtract or can you do any ?

OpenStudy (anonymous):

ur always adding not substracting

OpenStudy (13ziar):

sure?

OpenStudy (anonymous):

yeh

OpenStudy (13ziar):

ook thanks for helping jizakallah

OpenStudy (13ziar):

:)

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

I think this guy explain it better than me.. here you go https://www.youtube.com/watch?v=EooANDoThys